OpenStudy (anonymous):

calculate the unknown side (x) of the triangle shown below. Give your answer to one (1) digit after the decimal.

OpenStudy (anonymous):



OpenStudy (anonymous):

If that angle is 90 degrees, then you should use the pythagorean theorem of right triangles to find x.

OpenStudy (anonymous):

Using the Pythagorean theorem in a right triangle, the hypotenuse squared is equal to the sum of the squares of the remaining sides. In your case, A=21, B=13, and C=x. Pythagorean Theorem states \[A^2+B^2=C^2\].

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

\[21^{2} + 13^{2} = z^{2}\] by pythagorean theorem

OpenStudy (anonymous):

oh okay guys got it

OpenStudy (anonymous):

x=610 wich is equal to 24.70 correct?

OpenStudy (tkhunny):

Did you just write 610 = 24.70? If you mean the Square Root of 610 is close to 24.70, that that is much more sensible.

OpenStudy (anonymous):

610 i s close to 24,70